Problem 17-3 Cardinal Paz Corp, carmies an account in its general ledger called Investments, which contained debits for investment purchases, and no credits, with the following descriptions Feb. 1, 2017 Sharapova Company common stock, $100 par, 200 shares April July 1 537,400 us government bonds, 11%, due April, 2027, nterest payable Aprn and October 1, 110 bonds of $1,000 par each 110,000 . McGrathCompany 12% bonds, par $50,000, dated March 1, 2017, purchased at 104 plus accrued interest, interest payable annually on March 1, due March 1, 2037 54,000 Your answer is partially correct. Try again. Prepare entries necessary to classify the amounts into proper accounts accounts, assuming that all the securities are -sale. (Credit account titles are automatically indented when amount is entered Do not indent manually. If no entry is muired, selectNo Entry-for the the amounts.) account titles and enter o for Debit Credit Account Titles and Explanation Your answer is partially correct. Try again. Prepare the entry to record the accrued interest and the amortization of premium on December 31, 2017, using the straight-line method. (Round answers to 0 decimal places, e.g. 2,500. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter O for the amounts.) Date Account Titles and Explanation Debit Credit Dec. 31, 2017 Interest Receivable Your answer is incorrect. Try again. The fair values of the investments on December 31, 2017, were: Sharapova Company common stock U.S. government bonds McGrath Company bonds $31,800 124,700 58,600 What entry, if any, would you recommend be made? (Round answers to 0 decimal places, eg 2soo. credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter 0 for the amounts.) Date Account Titles and Explanation Debit Credit Dec. 31, 2017 (Entry for debt investment)
Answer #1

Part 1

Particulars Debit Credit
Equity Investment (Available for sale) $    37,400.00
Debt Investment (Available for sale) $ 162,000.00
[$110,000 + 104%($50,000)]
Interest Revenue $      2,000.00
($50,000 X 12%) x 4/12
Investment $ 201,400.00

[note - Above entry can be passed by adding equity investment and debt investment together under Available for sale A/C.
i.e the alternative entry can be -
Available for sale Dr $199400 ,
Interest revenue Dr $2000,
Investment Cr $201400]

Part - 2

Company bond due date = 2037, 1st march
Company bond purchase date = 2017, 1st july
Holding period = 19 years 8 month = 236 months
Premium Amortisation = [$2000 x 6/236] = $51
Accrued Interest on govt. bond =( $110,000 x 11%) x 3/12 = $3025
Accrued Interest on company bond = ($50,000 x 12%) x 10/12 = $5000

Interest Revenue = $(3025 + 5000 - 51) = $7974

Date Particulars Debit Credit
31-Dec-17 Interest receivables $ 8,025.00
Interest revenue $    7,974.00
[See working note]
Debt Investment (Available for sale) $          51.00
[See working note]

Part 3

Securities available for sale Cost Fair Value Unrealised gain (loss)
Common Stock $    37,400.00 $    31,800.00 $                      (5,600.00)
Govt Bond $ 110,000.00 $ 124,700.00 $                      14,700.00
Company bond $    51,949.00 $    58,600.00 $                        6,651.00
($52000 - $51) $                                     -  
Total $                      15,751.00
Particulars Debit Credit
Available for Sales (Fair value Adjustments) $ 15,751.00
Unrealised Gain $ 15,751.00

[Note - Available for sales A/C can be shown in Equity invetment 5600 and Debt Investment $(14700 + + 6651) = $21351
i.e. An alternative entry can be -
Debt Investment (Available for sales) Dr $21351
Equity Investment (Available for sales) Cr $5600
Unrealised Gain $15751]